Question 4

Orville Company’s standard and actual costs per unit are provided below for the most recent period. During this time period 400 units were actually produced.

Standard

Actual

Materials:

     Standard: 2 metres at $1.50 per m.

$3.00

     Actual: 2.1 metres at $1.60 per m.

$3.36

Direct labour:

     Standard: 1.5 hrs. at $6.00 per hr.

9.00

     Actual: 1.4 hrs. at $6.50 per hr.

9.10

Variable overhead:

     Standard: 1.5 hrs. at $3.40 per hr.

5.10

     Actual: 1.4 hrs. at $3.10 per hr.

4.34

Total unit cost

$17.19

$16.80

For simplicity, assume there was no inventory of materials at the beginning or end of the period.

Required:

Given the information above, compute the following variances. Also indicate if the variances are favorable or unfavorable.

  1. Materials price variance
  2. Materials quantity variance
  3. Direct labour rate variance
  4. Direct labour efficiency variance
  5. Variable overhead spending variance
  6. Variable overhead efficiency variance
Answer #1

Materials price variance = AQ*(AP-SP) = (400*2.1) *(1.60-1.50) = $84 Unfavorable

Material quantity variance= SP*(AQ-SQ) =1.50*(840-800) =$60 unfavorable

Direct labor rate variance= AH*(AR-SR) = 560*(6.50-6.00) = $280 unfavorable

Direct labor efficiency variance= SR*(AH-SH) = 6.00*((400*1.4) -(400*1.5)) = -240= $240 Favorable

Variable overhead spending variance= AH*(AR-SR) = 560*(3.10-3.40)) = =-168 =$168 favorable

Variable overhead efficiency variance= SR*(AH-SH) = 3.40*(560-600) = -136 = $136 favorable